OpenStudy (anonymous):

HOW MANY solutions does the equation have ? 2(2x+5)=4(x+3)

OpenStudy (anonymous):

Well. Lets solve it first. Distribute both sides.

OpenStudy (anonymous):

Are you not going to reply to your own question...?

OpenStudy (anonymous):

2(2x+5)=4(x+3) Distribute 4x+10=4x+12